40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog, Page 5 of 9  Not logged in ELOG logo
ID Date Author Type Category Subject
  211   Sun Jul 20 17:19:50 2014 KojiMechanicsCharacterizationI1OMC vibration test ~ 2nd round

Improved vibration measurement of the OMC

Improvement

- Added some vibration isolation. Four 1/2" rubber legs were added between the OMC bread board and the transport fixture (via Al foils).
  In order to keep the beam height same, 1/2" pedestal legs were removed.

- The HEPA filter at the OMC side was stopped to reduce the excitation of the breadboard. It was confirmed that the particle level for 0.3um
  was still zero only with the other HEPA filter.


Method

- Same measurement method as the previous entry was used.

Results

Breadboard

- In this new setup, we could expect that the resonant frequency of the body modes were close to the free resonances, and thus the Q is higher.
  Noise is much more reduced and it is clear that the resonance seen 1.1kHz is definitely associated with the body mode of the breadboard (red curve).

  As a confirmation, some metal objects were placed on the breadboard as tried before. This indeed reduced the resonant frequency (blue curve).

I1OMC_vibration_test_Breadboard.pngI1OMC_vibration_test_Breadboard_HiRes.png

DCPD / QPD

- Vibration on the DCPDs and QPDs mainly excited the modes above 2~3kHz.
  In order to check if they are coming from the housing, we should run FEA models.

- Some excitation of the breadboard mode at 1.1kHz was also seen.

I1OMC_vibration_test_DCPD.pngI1OMC_vibration_test_QPD.png

CM1/CM2 (PZT mirrors)

- Baseically excitation was dominated by the PZT mode at 10kHz. Some spourious resonances are seen at 4~5kHz but I believe this is associated with the weight placed on the excitation PZT.

I1OMC_vibration_test_CM.png

FM1/FM2 and peripheral prism mirrors (BSs and SMs)

- The modes of the FMs are seen ~8k or 12kHz. I believe they are lowered by the weight for the measurement. In any case, the mode frequency is quite high compared to our frequency region of interest.

- As the prism resonance is quite high, the excitation is directly transmitted to the breadboard. Therefore the excitation of the non-cavity caused similar effect to the excitation on the breadboard.
  In fact what we can see from the plot is excitation of the 1.1kHz body mode and many high frequency resonances.

I1OMC_vibration_test_FM.pngI1OMC_vibration_test_Prism.png

Beam dumps

- This is also similar to the case of the peripheral mirrors.

I1OMC_vibration_test_BD.png

Attachment 1: I1OMC_vibration_test.pdf
I1OMC_vibration_test.pdf I1OMC_vibration_test.pdf I1OMC_vibration_test.pdf I1OMC_vibration_test.pdf I1OMC_vibration_test.pdf I1OMC_vibration_test.pdf I1OMC_vibration_test.pdf I1OMC_vibration_test.pdf
  210   Thu Jul 17 02:19:20 2014 KojiMechanicsCharacterizationI1OMC vibration test

Summary

- The breadboard has a resonance at 1.2kHz. The resonant freq may be chagned depending on the additional mass and the boundary condition.

- There is no forest of resonances at around 1kHz. A couple of resonances It mainly starts at 5kHz.

- The PZT mirrors (CM1/CM2) have the resonance at 10kHz as I saw in the past PZT test.


Motivation

- Zach's LLO OMC characterization revealed that the OMC length signals have forest of spikes at 400-500Hz and 1kHz regions.

- He tried to excite these peaks assuming they were coming from mechanical systems. It was hard to excite with the OMC PZT,
but actuating the OMCS slightly excited them. (This entry)

Because the OMC length control loop can't suppress these peaks due to their high frequency and high amplitude, they limit
the OMC residual RMS motion. This may cause the coupling of the OMC length noise into the intensity of the transmitted light.
We want to eventually suppress or eliminate these peaks.

By this vibration test we want to:

- confirm whether the peaks are coming from the OMC or not.
- identify what is causing the peaks if they are originated from the OMC
- correct experimental data for comparison with FEA

Method

- Place a NOLIAC PZT on the object to be excited.
- Look at the actuation signal for the OMC locking to find the excited peaks.

Results

Breadboard

- This configuration excited the modes between 800-1.2kHz most (red curve). As well as the others, the structures above 5kHz are also excited.

- The mode at 1.2kHz was suspected to be the bending mode of the breadboard. To confirm it, metal blocks (QPD housing and a 4" pedestal rod)
  were added on the breadboard to change the load. This actually moved (or damped) the mode (red curve).

- Note that the four corners of the breadboard were held with a PEEK pieces on the transport fixture.
  In addition, the installed OMC has additional counter balance mass on it.
  This means that the actual resonant frequency can be different from the one seen in this experiment. This should be confirmed with an FEA model.
  The breadboard should also exhibit higher Q on the OMCS due to its cleaner boundary condition. 

 

I1OMC_vibration_test_Breadboard.png

DCPD / QPD

- Vibration on the DCPDs and QPDs mainly excited the modes above 3kHz. The resonances between 3 to 5kHz are observed in addition to the ubiquitous peaks above 5kHz.
  So are these coming from the housing? This also can be confirmed with an FEA model.

- Some excitation of the breadboard mode at 1.2kHz is also seen.

 

I1OMC_vibration_test_DCPD.pngI1OMC_vibration_test_QPD.png

CM1/CM2 (PZT mirrors)

- It is very obvious that there is a resonance at 10kHz. This was also seen in the past PZT test. This can be concluded that the serial resonance of the PZT and the curved mirror.
- There is another unknown mode at around 5~6kHz.

- Some excitation of the breadboard mode at 1.2kHz is also seen.

I1OMC_vibration_test_CM.png

FM1/FM2 and Peripheral prism mirrors (BSs and SMs)

- They are all prism mirrors with the same bonding method.

- The excitation is concentrated above 5kHz. Small excitation of the breadboard mode at 1.2kHz is also seen. Some bump ~1.4kHz is also seen in some cases.

I1OMC_vibration_test_FM.png I1OMC_vibration_test_Prism.png

Beam dumps

- The excitation is quite similar to the case of the peripheral mirrors. Some bump at 1.3kHz.

I1OMC_vibration_test_BD.png


Other tapping test of the non-OMC object on the table

- Transport fixture: long side 700Hz, short side 3k. This 3K is often seen in the above PZT excitation

- Fiber coupler: 200Hz and 350Hz.

- The beam splitter for the back scattering test: 900Hz

  209   Tue Jul 15 03:34:16 2014 KojiOpticsCharacterizationOMC backscatter measurement

Backscatter measurement ~ 2nd round


Summary

- The backscatter reflectivity of the 3rd OMC is 0.71 ppm

- From the spacial power distribution, it is likely that this is not the upper limit but the actual specular spot from the OMC,
propagating back through the input path.


Improvement

- The power meter was heavily baffled with anodized Al plates and Al foils. This reduced many spourious contributions from the REFL path and the input beam path.
  Basically, the power meter should not see any high power path.

- The beam dump for the forward going beam, the beamsplitter, and the mirrors on the periscope were cleaned.

- The power meter is now farther back from the BS to reduce the exposed solid angle to the diffused light

- The REFL path was rebuilt so that the solid angle of the PD was reduced.

OMC_backscatter.png


Backscattering measurement

- Pin = 12.3 +/- 0.001 [mW]

- RBS = 0.549 +/- 0.005

- Pback = 4.8 +/- 0.05 [nW] (OMC locked)       ==> ROMC(LOCKED) = 0.71 +/- 0.01 [ppm]

- Pback = 3.9 +/- 0.05 [nW] (OMC unlocked)   ==> ROMC(UNLOCKED) = 0.57 +/- 0.01 [ppm]

Note that the aperture size of Iris(B) was ~5.5mm in diameter. 


V-dump test

- Additional beam dump (CLASS A) was brought from the 40m. This allowed us to use the beam dump before and after the periscope.

- When the beam dump was placed after the periscope: P = 0.9+/-0.05nW

- When the beam dump was placed before the periscope: P=1.0+/-0.1nW

===> This basically suggests that the periscope mirrors have no contribution to the reflected power.

- When the beam dump was placed in the REFL path: P=2.1+/-0.1nW


Trial to find backward circulating beam at the output coupler

The same amount of backreflection beam can be found not only at the input side of the OMC but also transmission side.
However, this beam is expected to be blocked by the beamsplitter. It was tried to insert a sensor card between the output coupler
and the transmission BS, but nothing was found.


In order to see if the detected power is diffused light or not, the dependence of the detected light power on the aperture size was measured.
Note that the dark offset was nulled during the measurement.

IRIS B
aperture   detected
diameter   power

[mm]       [nW]
 1.0        1.1

 2.5        2.6
 4.25       4.0
 5.5        4.6
 8.0        5.3
 9.0        6.1
11.0        6.3
15.0        7.0

We can convert these numbers to calculate the power density in the each ring. 
(Differentiate the detected power and aperture area. Calculate the power density in each ring section, and plot them as a function of the aperture radius)


This means that the detected power is concentrated at the central area of the aperture.
(Note that the vertical axis is logarithmic)

If the detected power is coming from a diffused beam, the power density should be uniform.
Therefore this result strongly suggests that the detected power is not a diffused beam but
a reflected beam from the OMC.

According to this result, the aperture size of 2.6mm in raduis (5.5mm in diameter) was determined for the final reflected power measurement.

Attachment 1: OMC_backscatter.pdf
OMC_backscatter.pdf
  208   Tue Jul 15 03:00:42 2014 KojiOpticsCharacterizationOMC backscatter measurement

Presence of the misaligned SRM (T=20%) was forgotten in the previous entry.
This effectively reduces the OMC reflectivity by factor of 25.

This is now reflected in the original entry. Also the argument about the power spectram density was modified.

Quote:

First, I'm looking at the alog by Zach: https://alog.ligo-la.caltech.edu/aLOG/index.php?callRep=8674

I'm not sure how this measurement can be converted into RIN. Well, let's try. Assuming his measurement is done with the single bounce beam from an ITM,
and assuming this plot is already normalized for RIN, we may need to multiply the number on the plot by factor of two or so. Then it's about factor of 5 lower RIN
than the expected RIN. And in terms of R, it is 25 times lower.

 

  207   Sun Jul 13 17:46:28 2014 KojiOpticsCharacterizationOMC backscatter measurement

Backscattering reflectivity of the 3rdOMC was measured.


Attached: Measurement setup

1) A CVI 45P 50:50 BS was inserted in the input beam path. This BS was tilted from the nominal 45 deg so that the reflection of the input beam is properly dumped.
This yielded the reflectivity of the BS deviated from 45deg. The measured BS reflectivity is 55%+/-1%.

2) The backward propagating beam was reflected by this BS. The reflected beam power was measured with a powermeter.

3) The powermeter was aligned with the beam retroreflected from the REFL PDH and the iris in the input path. The iris was removed during the measurement
as it causes a significant scatter during the measurement.

4) While the cavity was either locked or unlocked, no visible spot was found at the powermeter side.


The input power to the OMC was 14.6mW. The detected power on the powermeter was 66.0+/-0.2nW and 73.4+/-0.3nW with the cavity locked and unlocked, respectively.
This number is obtained after subtraction of the dark offset of 5.4nW.

Considering the reflectivity of the BS (55+/-1%) , the upper limit of the OMC reflectivity (in power) is 8.18+/-0.08ppm and 9.09+/-0.09ppm for the OMC locked and unlocked respectively. Note that this suggests that the REFL path has worse scattering than the OMC cavity but it is not a enough information to separate each contribution to the total amount.


Impact on the OMC transmission RIN in aLIGO:

- The obtained reflectivity (in power) was 8ppm.
- For now, let's suppose all of this detected beam power has the correct mode for the IFO.
- If the isolation of the output faraday as 30dB is considered, R=8e-9 in power reaches the IFO.
- The IFO is rather low loss when it is seen as a high reflector from the AS port.
- Thus this is the amount of the light power which couples to the main carrier beam.

When the phase of the backscattered electric field varies, PM and AM are produced. Here the AM cause
the noise in DC readout. Particularly, this recombination phase is changing more than 2 pi, the fringing
between the main carrier and the backscattered field causes the AM with RIN of 2 Sqrt(R).

Therefore, RIN ~ 2e-4 is expected from the above of backscattering.


Now I'm looking for some measurement to be compared to with this number.

First, I'm looking at the alog by Zach: https://alog.ligo-la.caltech.edu/aLOG/index.php?callRep=8674

I'm not sure how this measurement can be converted into RIN. Well, let's try. Zach told me that the measured value is already normalized to RIN.
He told me that the modulation was applied at around 0.1Hz. The maximum fringe velocity was 150Hz from the plot.
At 100Hz, let's say, the RIN is 2e-6 /rtHz. The fringe speed at 100Hz is ~70Hz/sec. Therefore the measurement stays in the 100Hz freq bin
only for delta_f/70 = 0.375/70 = 5.3e-3 second. This reduces the power in the bin by sqrt(5.3e-3) = 0.073.

2e-6 = 2 sqrt(R) *0.73 ==> R = 2e-10

This number is for the combined reflectivity of the OMC and the OMC path. Assuming 30dB isolation of the output Faraday
and 20% transmission of SRM, the OMC reflectivity was 5e-6. This is in fact similar number to the measured value.

If I look at the OMC design document (T1000276, P.4), it mentions the calculated OMC reflection by Peter and the eLIGO measurement by Valera.
They suggests the power reflectivity of the order of 1e-8 or 1e-7 in the worst case. This should be compared to 8ppm.
So it seems that my measurement is way too high to say anything useful. Or in the worst case it creates a disastrous backscattering noise.


So, how can I make the measurement improved by factor of 100 (in power)

- Confirm if the scattering is coming from the OMC or something else. Place a good beam dump right before the OMC?

- Should I put an aperture right before the power meter to lmit the diffused (ambient) scatter coming into the detector?
  For the same purpose, should I cover the input optics with an Al foil?

- Is the powermeter not suitable for this purpose? Should I use a PD and a chopper in front of the OMC?
  It is quite tight in terms of the space though.

- Any other possibility?

Attachment 1: OMC_backscatter.pdf
OMC_backscatter.pdf
  206   Fri Jul 11 00:06:33 2014 KojiOpticsCharacterizationI1OMC PD

DCPD#             DCPD1      DCPD2
Housing#          #009       #010
Diode#            #07        #10
Shim              1.00mm 01  1.00mm 02   (1.00mm = D1201467-09)

-------------------------------------
Power Incident     11.1 mW   10.6 mW
Vout                7.65 V    7.33 V

Responsivity[A/W]   0.69      0.69
Q.E.                0.80      0.81
-------------------------------------
photo              2nd        1st

 

PD alignment confirmation

  205   Thu Jul 10 23:22:28 2014 KojiOpticsCharacterizationI1OMC QPD

QPD#              QPD1       QPD2
Housing#          #006       #007
Diode#            #50        #51
Shim              1.25mm 03  1.25mm 02   (1.25mm = D1201467-10)

-------------------------------------
Power Incident    123.1-13.0 uW  124.5-8.0 uW
Sum Out            77.0 mV   82.5 mV
Vertical Out      -24.0 mV  - 8.8 mV
Horizontal Out      4.2 mV    9.0 mV
SEG1              -11.6 mV  -16.0 mV
SEG2              -12.6 mV  -18.0 mV
SEG3              -25.2 mV  -24.4 mV
SEG4              -21.4 mV  -21.4 mV
-------------------------------------
Spot position X   -21   um  -19   um  (positive = more power on SEG1 and SEG4)
Spot position Y   +102  um  +47   um  (positive = more power on SEG3 and SEG4)
-------------------------------------

Responsivity[A/W] 0.70      0.71
Q.E.              0.82      0.83
-------------------------------------

Arrangement of the segments
View from the beam
/ 2 | 1 X
|---+---|
\ 3 | 4 /

---------------

I(w,x,y) = Exp[-2 (x^2 + y^2)/w^2]/(Pi w^2/2)

(SEG_A+SEG_B-SEG_C-SEG_D)/(SEG_A+SEG_B+SEG_C+SEG_D) = Erf[sqrt(2) d/w]

d: distance of the spot from the center
w: beam width

  204   Thu Jul 10 08:34:57 2014 KojiGeneralGeneralTo Do

Optical tests

  • Cleaning
  • Power Budget
  • FSR measurement
  • TMS measurement
  • TMS measurement (with DC voltage on PZTs)
  • PZT DC response
  • PZT AC response
  • QPD alignment
  • DCPD alignment

Backscattering test

Cabling / Wiring

  • Attaching cable/mass platforms
  • PZT cabling
  • DCPD cabling (to be done at LHO)
  • QPD cabling (to be done at LHO)

Vibration test

Baking

First Contact

Packing / Shipping

  203   Thu Jul 10 01:39:38 2014 KojiElectronicsGeneralPZT wire

Rich came to the OMC lab. Pins for the mighty mouse connector were crimped on the 4 PZT wires.
We found the male 4pin mighty mouse connector in the C&B area.

The cable inventory was checked with ICS/DCC combo. It turned out that most of the on-board cables
are at LHO. We decided to send the OMC there and then the cables are installed at the site.

Attachment 1: P7096669.JPG
P7096669.JPG
  202   Tue Jul 8 18:54:54 2014 KojiMechanicsCharacterizationPZT characterization

Each PZT was swept with 0-150V 11Hz triangular wave.
Time series data for 0.2sec was recorded for each PZT.

The swept voltage at the resonances were extracted and the fringe number was counted.
Some hysteresis is seen as usual.

The upward/downward slopes are fitted by a linear line.

The average displacement is 11.3nm/V for PZT1 and 12.7nm/V.

The PZT response was measured with a FFT analyzer. The DC calibration was adjusted by the above numbers.

Attachment 1: PZT_Scan.pdf
PZT_Scan.pdf
Attachment 2: I1OMC_PZT_Response.pdf
I1OMC_PZT_Response.pdf
  201   Tue Jul 8 04:08:06 2014 KojiGeneralGeneralExpoxy reapplication for beam dumps

Firstly, the excess epoxy was removed using a cleaned razor balde

Secondly, EP30-2 epoxy was applied at the exterior edges of the beam dump.
Interior of the V were glued at two points. This is to keep the gap away from being trapped

Here is the result of the gluing. Some epoxy was sucked into the gap by capillary action.
I believe, most of the rigidity is proivded by the bonds at the edges.

  200   Mon Jul 7 01:36:03 2014 KojiGeneralGeneralTo Do

Optical tests

  • Cleaning
  • Power Budget
  • FSR measurement
  • TMS measurement
  • TMS measurement (with DC voltage on PZTs)
  • PZT DC response
  • PZT AC response
  • QPD alignment
  • DCPD alignment

Backscattering test

Cabling / Wiring

  • Attaching cable/mass platforms
  • PZT cabling
  • DCPD cabling
  • QPD cabling

Vibration test

Baking

First Contact

Packing / Shipping

  199   Sun Jul 6 08:31:14 2014 KojiOpticsCharacterizationHOM measurement with PZT vol swept

3rd OMC, HOM diagram at PZT1=0V and PZT2=50V.

First coincidence with the carrier is the 32nd-order carrier mode. Very good.

Attachment 1: HOM_plot.pdf
HOM_plot.pdf
  198   Sun Jul 6 03:56:40 2014 KojiOpticsCharacterizationHOM measurement with PZT vol swept

Cavity FSR/TMS measurement (2014/7/5) with PZT voltages swept from 0V to 200V (50V step)

Attachment 1: Cav_scan_response_PZT1.pdf
Cav_scan_response_PZT1.pdf Cav_scan_response_PZT1.pdf Cav_scan_response_PZT1.pdf Cav_scan_response_PZT1.pdf Cav_scan_response_PZT1.pdf Cav_scan_response_PZT1.pdf Cav_scan_response_PZT1.pdf Cav_scan_response_PZT1.pdf
Attachment 2: Cav_scan_response_PZT2.pdf
Cav_scan_response_PZT2.pdf Cav_scan_response_PZT2.pdf Cav_scan_response_PZT2.pdf Cav_scan_response_PZT2.pdf Cav_scan_response_PZT2.pdf Cav_scan_response_PZT2.pdf Cav_scan_response_PZT2.pdf Cav_scan_response_PZT2.pdf
Attachment 3: OMC_HOM_140705.pdf
OMC_HOM_140705.pdf
  197   Sun Jul 6 02:46:20 2014 KojiOpticsCharacterizationOMC power budget

3rd OMC power budget (2014/7/2)

Input power: 34.8mW

REFLPD dark offset:  -7.57mV
REFLPD unlocked: 6.22 V
REFLPD locked: 110mV

Transmitted Power: 16.8mW (T) and 15.9mW (R)
CM1 transmission: 0.176mW
CM2 transmission: 0.181mW

Cavity Finesse: 399.73


Junk light: 0.64mW (out of 34.8mW)
Coupled beam: 34.16 mW (out of 34.8mW)
Mode Matching: 0.982
Cavity reflectivity: 467ppm
Loss per mirror in ppm: 63.8ppm
Cavity transmission (for TEM00 carrier): 0.957

FM1: R = 0.992277, T = 7659.46
FM2: R = 0.992277, T = 7659.46
CM1: R = 0.999895, T = 41.5461
CM2: R = 0.999893, T = 42.7309


Compare the above number with the best result obtained during the alignment trials

Input power: 34.4mW

REFLPD dark offset:  -7.5mV
REFLPD unlocked: 5.99 V
REFLPD locked: 104mV

Transmitted Power: Total 32.7mW (T+R)
CM1 transmission: 0.194mW
CM2 transmission: 0.194mW

Cavity Finesse: 400


Junk light: 0.631mW (out of 34.4mW)
Coupled beam: 33.77 mW (out of 34.4mW)
Mode Matching: 0.982
Cavity reflectivity: 255ppm
Loss per mirror in ppm: 39.7ppm
Cavity transmission (for TEM00 carrier): 0.968


  196   Sun Jul 6 02:45:56 2014 KojiOpticsGeneralFSR Measurement

3rd OMC FSR / Finesse measurement

RF AM was injected by detuning a HWP.

Attachment 1: finesse_measurements_log.pdf
finesse_measurements_log.pdf
  195   Thu Jul 3 17:45:18 2014 KojiGeneralGeneralBeam dump delamination

Here is the resolution.

I'll apply fillets of EP30-2 along the edges of the black glass (See figure).
In order to allow the air escape from the gap, the inside of the V will not be painted.
In any case, I don't have a good access to the interior of the V.

Dennis assured that the outgassing level will be ok even if the EP30-2 is cured at the room temp if the mixture is good.
But just in case, we should run an RGA scan (after 50degC for 24hour vac bake).
I prefer to do this RGA scan right after all of the test and cabling and right before the shipment.
Dennis is checking if we can even waive the RGA scan owing to the small volume of the glue.

beamdump_delamination_solution.png

  194   Wed Jul 2 18:58:42 2014 KojiGeneralGeneralBeam dump delamination

While the OMC breadboard was being inspected, it was found that two out of five black-glass beam dumps showed sign of delamination.
(attached photos).

The base of the each beam dump is a fused silica disk (25mm dia.). The black glass pieces are bonded to the disk. The bond is EP30-2
epoxy without glass beads for bond lining. The disk is bonded on the fused silica bread board with Optocast UV low-viscous epoxy.
The delamination is about 70% of the bonded area. They don't seem to fall off immediately. But the glass pieces are not completely secure.
(i.e. finger touch can change the newton ring fringes) So there might be some risk of falling off during transportation.

The engineering team and I are exploring the way to secure them in-situ, including the method to apply UV epoxy with capillary action.

Attachment 1: beamdump_delamination.png
beamdump_delamination.png
  193   Wed Jul 2 16:41:43 2014 KojiGeneralGeneralOMC baking

OMC is back from the oven today.

To Do:

Optical tests

  • Cleaning
  • Power Budget
  • FSR measurement
  • TMS measurement
  • TMS measurement (with DC voltage on PZTs)
  • PZT DC response
  • PZT AC response
  • QPD alignment
  • DCPD alignment
  • First Contact

Backscattering test

Cabling / Wiring

  • Attaching cable/mass platforms
  • PZT cabling
  • DCPD cabling
  • QPD cabling

Vibration test

Packing / Shipping

  192   Fri Jun 27 18:51:33 2014 KojiGeneralGeneralSupply

PTOUCH TAPE (12mm white) x 2

9V batteries

  191   Fri Jun 27 12:29:50 2014 KojiGeneralGeneralOMC baking

The OMC went into the oven at around 2PM on Thursday. It will be baked at 80degC for 48 hours.
The RGA result will be obtained on Monday.

Link to the ICS entry

P6266536.jpg

  190   Tue Jun 24 18:43:15 2014 KojiGeneralGeneralICS entries for the OMC baking/assembly

For baking:

  • Assembly Name aLIGO Output Mode Cleaner
    Assembly Number D1201439
     
  • Part Name: Breadboard transport fixture
    Part # / Drawing #: D1201515
     

TO BE ADDED TO THE ASSEMBLY after the bake: [DONE]
803-003-07M6-4PN-598A-0-Bulk-H42Q001
D1201274-V1-00-S009: OMC DCPD Housing        (remove part)
D1201274-V1-00-S010: OMC DCPD Housing        (remove part)
D1201275-V1-00-0006: OMC DCPD FACE PLATE        (remove part)
D1201275-V1-00-0007: OMC DCPD FACE PLATE        (remove part)
D1201280-V1-00-0006: OMC QPD HOUSING        (remove part)
D1201280-V1-00-0007: OMC QPD HOUSING        (remove part)
D1201281-V1-00-0006: OMC QPD FACE PLATE        (remove part)
D1201281-V1-00-0007: OMC QPD FACE PLATE        (remove part)
D1300052-V1-00-0003: aLIGO OMC BRACKET, CABLE CONNECTOR        (remove part)
D1300057-v2-00-0021: aLIGO CABLE PEG        (remove part)
D1300057-v2-00-0022: aLIGO CABLE PEG        (remove part)
D1300057-v2-00-0023: aLIGO CABLE PEG        (remove part)
D1300057-v2-00-0024: aLIGO CABLE PEG        (remove part)
D1300057-v2-00-0025: aLIGO CABLE PEG        (remove part)
D1300057-v2-00-0026: aLIGO CABLE PEG        (remove part)
D1300057-v2-00-0027: aLIGO CABLE PEG        (remove part)
D1300057-v2-00-0028: aLIGO CABLE PEG        (remove part)
D1300057-v2-00-0029: aLIGO CABLE PEG        (remove part)
D1300057-v2-00-0030: aLIGO CABLE PEG        (remove part)
D1300060-V1-00-0005: aLIGO OMC BRACKET, MASS MOUNTING        (remove part)
D1300060-V1-00-0006: aLIGO OMC BRACKET, MASS MOUNTING        (remove part)


====================================
More entries to be added (Found in the LHO OMC entry) [DONE]
D1300371-V2-00-S1301806: ISC DCPD Cable for OMC-Breadboard Bracket to DCPD #1
D1300372-V2-00-S1301807: ISC DCPD Cable for OMC-Breadboard Bracket to DCPD #2
D1300373-V3-00-S1301810: ISC QPD Cable for OMC-Structure to Breadboard Bracket
D1300374-V2-00-S1301813: ISC QPD Cable for OMC-Breadboard Bracket to QPD #1

  189   Mon Jun 23 21:54:16 2014 KojiOpticsGeneralAll of the gluing completed

The bottom-side templates were removed.

The last beam dump was removed

TODO

ICS entry

Bring the OMC to the bake lab

Vacuum baking

Bring it back to the OMC lab

Cabling / Wiring

VIbratin test

Optical tests

Backscattering test

Packing / Shipping

 

  188   Fri Jun 20 18:59:12 2014 KojiOpticsGeneralAll of the invar blocks have been glued

All of the INVAR blocks have been glued.

I found thinner shims in the stock.

On Monday, the template will be removed.

EP30-2 7g mixed with 0.35g of 75-90um sphere


TODO

EP30-2 gluing of the INVAR blocks for the PDs

PDs/QPDs need to be slightly lower -> order more shims

Remove the templates

Glue the last beam dump

Vibration test?

Bring the OMC to the bake lab

Vacuum baking

Bring it back to the OMC lab

Cabling / Wiring

Optical tests

Backscattering test

Packing / Shipping

  187   Thu Jun 19 23:16:50 2014 KojiOpticsGeneralAll of the prisms have been glued

- All of the PRISM mirrors have been glued

- 4 out of 5 beam dumps have been glued


TODO

EP30-2 gluing of the INVAR blocks for the PDs

PDs/QPDs need to be slightly lower -> order more shims

Remove the templates

Glue the last beam dump

Vibration test?

Bring the OMC to the bake lab

Vacuum baking

Bring it back to the OMC lab

Cabling / Wiring

Optical tests

Backscattering test

Packing / Shipping

P6196525.JPG

  186   Sat May 17 07:40:14 2014 KojiOpticsCharacterizationI1OMC cavity mirrors glued

I1OMC cavity mirrors were glued.

FSR = 264.82MHz => Lcav = 1.132m (nominal 1.132m)

TMS/FSR for Vpzt1=Vpzt2=0: 0.2185 (V) and 0.2196 (H) (nominal 0.219)

 

aLIGO OMC: Power Budget 2014/5/16

<<<Measured Values>>>
Input Power: 35.7 [mW]
Transmitted Power through FM2: 33.5 [mW]
Transmitted Power through CM1: 0.188 [mW]
Transmitted Power through CM2: 0.192 [mW]
Reflection PD DC output (Unlocked): 6.2 [V]
Reflection PD DC output (Locked): 0.096 [V]
Reflection PD DC output (Dark Offset): -0.00745 [V]
Assumed cavity finesse : 400.

<<<Results>>>
Input Power: 35.7 [mW]
Uncoupled light Power (Junk light + sidebands): 0.575698 [mW]
Input TEM00 Carrier Power: 35.1243 [mW]  (Ratio: 0.983874)
Cavity reflectivity (in power): 548.319 ppm
Cavity transmission (in power): 0.953756
Loss per mirror: 70.1183 ppm
FM1 power transmission: 7640.17 ppm
FM2 power transmission: 7640.17 ppm
CM1 power transmission: 43.2093 ppm
CM2 power transmission: 44.1337 ppm

 

Attachment 1: Cav_scan_response_140516_Pitch.pdf
Cav_scan_response_140516_Pitch.pdf
Attachment 2: Cav_scan_response_140516_Yaw.pdf
Cav_scan_response_140516_Yaw.pdf
  185   Fri May 16 00:13:36 2014 KojiOpticsCharacterizationCavity mirror gluing part 1

BS1/FM1/FM2 for I1OMC were glued.

FM1 had to be intentionally rotated.
FM1 had to be intentionally shifted to avoid scattering spot.

Pin: 36.3 / Ptrans: 33.7 = Raw transmission 92.8%
Vunlock = 6.30 / Vlock = 0.120

Mode matching (estim) 0.98
Loss per mirror 84ppm
Cavity transmission 0.947

ummm

Tomorrow:
- Transmission needs to be optimized
- Apply 50V to a PZT
- Cavity FSR/HOM should be optimized
- gluing

Put a cover
Return power meter / DC supply

  184   Wed May 14 02:15:15 2014 KojiOpticsCharacterizationFSR/TSM adjustment of the OMC cavity

1. FSR was adjusted and measured with "the golden arches" technique again.

FSR = 264.8412 MHz +/- 1400Hz => Lcav = 1.13197 m. (nominal 1.132m)

2. Transverse mode spacings for the vertical and horizontal modes were measured.

TMS/FSR = 0.218144 (V) / 0.219748 (H)

This is almost perfect!

The 19th-order lower sideband hit the resonance. Next step is to glue some of the flat mirrors.

Attachment 1: Cav_scan_response_140503_Pitch.pdf
Cav_scan_response_140503_Pitch.pdf
Attachment 2: Cav_scan_response_140503_Yaw.pdf
Cav_scan_response_140503_Yaw.pdf
  183   Mon May 12 22:43:02 2014 KojiOpticsCharacterizationMeasured FSR/TSM of the OMC cavity

Data analysis of the FSR/TSM measruement last week.

1. FSR was measured with "the golden arches" technique.

FSR = 263.0686 MHz +/- 900Hz

Lcav = 1.1396 m --> 7.6 mm too long! (nominal 1.132m)

2. Transverse mode spacings for the vertical and horizontal modes were measured.

TMS/FSR = 0.219366 (V) / 0.220230 (H) (Predicted value with the current cavity length 0.2196/0.2202 very close!)

We want to make this to be ~0.219 (~3% less)

With the current parameters, the 19th-order lower sideband make the coincident resonance.

Attachment 1: Cav_scan_response_Pitch.pdf
Cav_scan_response_Pitch.pdf
Attachment 2: Cav_scan_response_Yaw.pdf
Cav_scan_response_Yaw.pdf
  182   Thu Apr 17 21:39:25 2014 KojiOpticsGeneralMore alignment

STORY:

- The cavity mirrors have scattering spots. The cavity alignment should have been scanned to find a cavity mode to have lowest loss possible.
  BTW, We only have horizontal dof for the alignment scan.

- After some struggle nice cavity mode was found. The cavity transmission was 96% for the ideally matched TEM00 carrier.

- It turned out that this imposed too much beam shift in the input beam (~2mm).

- This big shift induces a lot of trouble for the peripheral optics (PDs, QPDs, sterring mirrors).

- What should we do???

Analysis:

- The beam needed to go up between CM1 and CM2 to have the right spots on them. ("UP" is the input side of the OMC).

- This imposed the beam between FM1 and FM2 moved up. In other word, for the given alignment of the FMs by the template,
  We needed to hit the upper part of the FMs to have the spots on the CMs up.

Solution:

- The above argument suggets that the nominal beam will give us the right spots on the CMs if we rotate the FMs.
  Of course this induces the spot move on the FMs. But this should not be the issue as the most of the loss seems to come from the CMs.

- How much misalignment show we give to the FMs? We want to shift the beam by 2mm on the CMs.
  The length of the optical lever is ~0.25m. Therefore the mialignment angle should be

  theta = 2e-3/2/0.25 = 4e-3 rad = 4mrad.

  The template pad has ~20mm separation. The thickness of the shim should be 20mm*4mrad = 80um

- Our aluminum foil seems to have the thickness of 30-40um. We can't have this minimum thickness on the template pad as there is not enough compression pressure
  => Just use a single layer of Al piece to shim the FMs.

Attempt:

- The shims were inserted at the upper pads of the FMs.

- Aligned the input beam and the CMs so that the spots on the CMs are approximately recovered.

- Measure the cavity power budget

Pin: 34.7mW
Refl PD: offset = -7.5mV, unlock = 6.07V, inlock = 89.7mV

Ptrans = 32.5mW

Ptrans(CM2) = 0.181mW
Ptrans(CM2) = 0.184mW

Assume finesse of 400

==>

Pin: 34.7mW
Pjunk: 0.534mW
Pcoupled: 34.1mW

Mode matching: 98.5%

Cavity reflectivity in power: 0.00061
Cavity transmission in power: 0.951 (This is not a best number but acceptable.)
Loss per mirror: 75.4ppm

FM power refl/trans: 0.9923 / 7630ppm
CM1 power refl/trans: 0.999882 / 42.8ppm
CM2 power refl/trans: 0.999881 / 43.5ppm
Total roundtrip loss of the cavity (Loss + CM leakage): 388ppm

Result:

How much the input beam is away from the left wall of the OMC breadboard?

40.88mm from the template edge
  8.36mm between the template edge and the bread board
=> 32.52mm

How much should this number be? 32.94mm from the solidworks model => With in 0.5mm! Nice!

Next:

- Just in case plce all of the optics and check if the beam is delivered within the alignment range of the optics

 

  181   Tue Mar 25 17:10:10 2014 KojiOpticsCharacterizationOMC spot position estimation

Spot positions were inferred from the photos

Attachment 1: OMC_spot.pdf
OMC_spot.pdf OMC_spot.pdf OMC_spot.pdf OMC_spot.pdf
  180   Mon Mar 3 02:46:21 2014 KojiGeneralCharacterizationSpot positions scanned

Spot positions on CM1 and CM2 scanned according to the recipes provided by the previous entry.

The best result obtained was:

Transmission from FM2: 32.7mW
Incident on BS1: 34.4mW

Reflection (Unlocked): 5.99V
Reflection (Locked): 104mV
Reflection (Dark): -7.5mV

to accomodate the spot on BS1 it had to be about a mm moved from the template.

This gives us:
- Portion of the TEM00 carrier: R = 1-(104+7.5)/(5990+7.5) = 0.981
- Raw transmission: 32.7/34.4 = 0.950
- TEM00 transmission 0.950/R = 0.969
- Excluding the transmission of BS1: 0.969/0.9926 = 0.976
  => loss per mirror ~40ppm

 

  179   Fri Feb 28 19:50:11 2014 KojiGeneralGeneralMisalignment ABCD matrix for the aLIGO OMC

Relationship between mirror misalignment and cavity mode shift was calculated.
The technique described in T0900647 by Sam Waldman was used.

The angles and displacement of the mirrors and beams are defined in the attached figure.

x1 = 0.893134 α + 1.10676 β + 1.32252 γ + 1.24619 δ
𝛳1 = 0.75864 α - 0.75864 β - 0.271075 γ + 0.271075 δ

x2 = 1.10676 α + 0.893134 β + 1.24619 γ + 1.32252 δ
𝛳2 = 0.75864 α + 1.24136 β - 0.271075 γ + 0.271075 δ

x3 = 1.32252 α + 1.24619 β + 1.1691 γ + 1.39962 δ
𝛳3 = -0.271075 α + 0.271075 β + 0.818668 γ - 0.818668 δ

x4 = 1.24619 α + 1.32252 β + 1.39962 γ + 1.1691 δ
𝛳4 = -1.24136 α - 0.75864 β - 0.271075 γ + 0.271075 δ

Assuming the flat mirrors are fixed:
If I want to move the x3 mirror up by 1mm without moving x4, the solution is
γ = -0.00197 mrad
δ = +0.00236 mrad

This yields:
x1 = +0.33mm, x2=+0.66mm, x3 = +1mm, x4 = 0mm

Attachment 1: misalignment.pdf
misalignment.pdf
  178   Tue Feb 18 18:58:38 2014 KojiGeneralGeneralLHO H1 OMC installation photos

LHO OMC installation photos

  177   Tue Dec 10 16:41:51 2013 KojiGeneralGeneralTo Buy

Masks / Wipes => done

  176   Thu Nov 21 00:05:35 2013 KojiGeneralGeneralH1OMC electronics arrangement

Checked the PZT arrangement: Mighty Mouse Pin1&2 -> PZT2 (DCPD side), Mighty Mouse Pin3&4 -> PZT1 (QPD side)

DCPD response:
Illuminate DCPD1 (T) -> DCPD B responded in MEDM
Illuminate DCPD2 (R) -> DCPD A responded in MEDM

QPD response:
Illuminate QPD1 -> QPD A responded in MEDM
Illuminate QPD2 -> QPD B responded in MEDM

--------

DCPD1 (T) is marked as "A". This PD is SN"0288"

DCPD2 (R) is marked as "B". This PD is SN"0721"

Corresponding iLOG for the performance

Attachment 1: H1OMC_cable_arrangement.pdf
H1OMC_cable_arrangement.pdf
  175   Mon Nov 4 19:43:56 2013 KojiGeneralGeneralH1OMC Packed

H1OMC PZT connector was replaced with the correct one. This was the final step for H1OMC.

Jeff and I packed the OMC and put it in the perikan case. It will be shipped tomorrow.

The other tools are also packed in the other box. Here is the list of the items

- Spare PD/QPDs (2 cages)
- Test PD/QPD cables
- Torque driver / bits
- Low noise transimpedance amp
- Kapton sheets
- First Contact kit
- 1/4-20 Screws for the balance weights
- OMC-Structure cables
- Preamp adapter plate
- Screws for the cable mounts
- Clean tools
  (scissors, tweezers, forceps, Diagonal pliers, long nose prier)
- Spare Peek cable ties
  174   Wed Oct 23 02:45:07 2013 KojiGeneralGeneralPD realignment

DCPD2 got misaligned during the cable installation. The PD alignment procedure have been gone through again.

Cavity locking

- Removed the FC layers for the cavity related mirrors.

- Aligned and locked the cavity.

PD alignment

- Loosen DCPD2. Checked the reflection with a IR card. Checked the spot on the PD with an IR viewer.

- Finger-tight the screws. Check the reflection with the card again. Check the pot on the PD with a CCD.

- If the spot positions are not satisfactory repeat the process.

- If the spot positions are satisfactory, take pictures of the CCD image.

- Fixing screws for all of the PDs/QPDs were tighten by the torque driver with a torque od 1.75 inch lb.

PD QE measurements

- Measure the power incident on the PDs.

- Set up the transimpedance amp to check the photo current.

- PD1 (T side) 9.10+/-0.03 V 13.02 +/- 0.01W -> QE ~80%

- PD2 (R side) 8.70+/-0.01 V 12.53 +/- 0.01W -> QE ~80%

- These are not strange values considering the presence of the glass caps.

PZT polarity check

- The connections between the PZT electrodes and the pins were checked.

- The positive side is marked by a knot on the wire.

FC painting

- The new FC bottle was brought from Downs, thanks to Margot.

 

  173   Tue Oct 22 17:17:59 2013 KojiGeneralGeneralH1OMC cabling

[Chub, Jeff, Koji]

We worked on the wiring and routing of the cables.

- The cables for the PZT was installed first.

- Pins for the mighty mouse connector were crimped on the PZT wires

- Checked the wiring diagram (D1300589) to find the pinouts.
  Pin1 of the mighty mouse is connected to PZT2+, Pin2 to PZT2-, PIn3 to PZT1+, and Pin4 to PZT1-

- Then QPD and PD cables are fixed on the cable harness.

- The QPD/PD cables are attached on the diode housings.

During this process one of the DCPD mounts moved. The fixing screws were not torqued enough.
This means that all of the FC layers need to be removed and the DCPD housing should be aligned again.

- We continued on the cabling. The cables were routed on the top (cable) side.

- Some of the cable pegs were tightened by PEEK cable ties.

- We found that Pin1 and Pin2 of the PZT cables were not intact anymore.

- We ask Chub to work on the PZT pins tomorrow. The PD alignment will be taken tonight or tomorrow.

  172   Wed Oct 16 19:16:29 2013 KojiOpticsCharacterizationPD alignment

 

 shim 1.5mm 001/002

  171   Tue Oct 15 18:50:08 2013 KojiOpticsCharacterizationQPD alignment

1) Deburr the bottom surfaces of the QPD housings

2) Aligned the QPDs

 

QPD#              QPD1       QPD2
Housing#          #004       #008
Diode#            #44        #46
Shim              1.75mm 001 1.25mm 001

-------------------------------------
Power Incident    125.7 uW  126.4 uW
Sum Out            80.1 mV   78.9 mV
Vertical Out      + 3.4 mV    0   mV
Horizontal Out    -23.7 mV  -26   mV
SEG1              -15.6 mV  -13.2 mV
SEG2              -13.1 mV  -13.3 mV
SEG3              -29.0 mV  -26.4 mV
SEG4              -23.2 mV  -26.3 mV
-------------------------------------
Spot position X   -13   um  - 0.8 um  (positive = more power on SEG1 and SEG4)
Spot position Y   +93   um +107   um  (positive = more power on SEG3 and SEG4)
-------------------------------------

Responsivity[A/W] 0.64      0.62
Q.E.              0.74      0.73
-------------------------------------

Arrangement of the segments
View from the beam
/ 2 | 1 X
|---+---|
\ 3 | 4 /

---------------

I(w,x,y) = Exp[-2 (x^2 + y^2)/w^2]/(Pi w^2/2)

(SEG_A+SEG_B-SEG_C-SEG_D)/(SEG_A+SEG_B+SEG_C+SEG_D) = Erf[sqrt(2) d/w]

d: distance of the spot from the center
w: beam width

  170   Mon Oct 14 15:50:55 2013 KojiOpticsCharacterizationH1 OMC Power budget

LHO OMC power budget

Date 2013/9/17 2013/9/17 2013/10/16 2013/10/22
Condition  Before the cleaning  After the cleaning  Confirmation  Confirmation
Input Power [mW]  35.2  35.4  34.54  34.9
REFLPD dark offset [V]  -0.00763  -0.00763  -0.00772  -0.000759
REFLPD unlocked [V]  0.0749 +/- 0.0005  0.067+/- 0.0005  0.0640+/-0.0005  0.0530+/-0.0001
REFLPD locked [V]  5.49 +/- 0.01  5.55+/-0.01  5.28+/-0.01  5.26+/-0.01
         
 Transmitted Power to DCPD1 (T) [mW]  16.5  16.4  16.1  16.0
 Transmitted Power to DCPD2 (R) [mW]  15.9  16.2  15.55  15.55
 FM2 transmission [mW]  32.4  32.9+/-0.1  -  -
 CM1 transmission [mW]  0.166  0.169  0.164  0.165
 CM2 transmission [mW]  0.165  0.169  0.158  0.162
 Input BS transmission [mW]  0.234  0.218  0.230  0.227
         
 Cavity Finesse  373.114  373.114  373.114  373.114
         
 Junk Light Power (Pjunk) [mW]  0.489  0.434  0.422  0.332
 Coupled beam power (Pcouple) [mW]  34.71  34.97  34.12  34.57
 Mode Matching (Pcouple/Pin) [mW]  0.986  0.988  0.988  0.990
 Cavity reflectivity in power  0.00115  0.00119  0.00136  0.00199
 Loss per mirror [ppm]  122  124  134  167
 Cavity transmission for TEM00 carrier
 0.933  0.932  0.927  0.913

 

Attachment 1: OMC_power_budget.pdf
OMC_power_budget.pdf OMC_power_budget.pdf OMC_power_budget.pdf OMC_power_budget.pdf
  169   Mon Oct 14 13:40:16 2013 KojiOpticsCharacterizationH1 OMC Optical testing

Since the middle of September, the optical tests of H1 OMC were took place.
Here is summary of the progress.

TEST1: FSR/FINESSE measurement before applying First Contact
TEST2: Power budget

MIrror cleaning with First Contact

TEST3: FSR/FINESSE measurement after First Contact application
TEST4: Power budget

TEST5: N/A

TEST6: HOM measurement @PZT V=0
TEST7: HOM measurement @PZT V=0-200

TEST8: DC response of the PZT
TEST9: AC response of the PZT

TEST10: PD/QPD alignment / output check

 

 

  168   Fri Sep 13 15:09:20 2013 KojiGeneralGeneralSprinkler installation: done

A sprinkler head was installed on the HEPA enclosure. The head is covered with a plastic cap.

Attachment 1: P9134379.jpg
P9134379.jpg
Attachment 2: P9134378.jpg
P9134378.jpg
  167   Sat Sep 7 17:20:56 2013 KojiGeneralGeneralOMC/PD lab optical table wrapping

[Koji Jeff]

In order to prepare for the splinkler installation on the HEPA enclosure, the table with the optics was wrapped with Ameristat sheets.

Attachment 1: P9064377.JPG
P9064377.JPG
  166   Wed Sep 4 22:22:54 2013 KojiGeneralGeneralH1 OMC wrapped and moved to the bake lab.

[Koji, Jeff]

We moved the H1OMC to the bake lab.

Chub set up the vacuum bake oven (Oven F) and running without the actual OMC.

We use low temperature (55degC) for the baking.

The actual OMC will be baked from tomorrow afternooon.

  165   Tue Sep 3 17:03:25 2013 KojiGeneralGeneralH1 OMC gluing completed

[Koji Jeff]

H1 OMC All Gluing completed

5 Glue H1 beam dumps (UV)

4 glass wire brackets glued on the H1 topside (UV) SN: #9/10/11/12

6 Invar blocks glued on the H1 topside (EP30) SN: #13/14/15/16/18/19

  164   Fri Aug 30 12:25:29 2013 KojiGeneralGeneralH1 OMC Invar mount gluing

The Invar Mounting Blocks were glued on the breadboard.

Serial number #1/2/5/6/7/8 -> I1 OMC cable side

Serial number #9/10/11/12 -> H1 OMC cavity side

Attachment 1: P8304368.JPG
P8304368.JPG
Attachment 2: P8304370.JPG
P8304370.JPG
  163   Fri Aug 30 12:24:28 2013 KojiOpticsCharacterizationH1OMC Spot positions

Beam heights on the diodes

DCPD1: 14.459mm -> With 1.5mm shim, the beam will be 0.038mm too low.

DCPD2: 14.221mm -> With 1.25mm shim, the beam will be 0.026mm too low.

QPD1: 14.691mm -> With 1.75mm shim, the beam will be 0.056mm too low.

QPD2: 14.379mm -> With 1.5mm shim, the beam will be 0.118mm too low.

Attachment 1: DCPD1.png
DCPD1.png
Attachment 2: DCPD2.png
DCPD2.png
Attachment 3: QPD1.png
QPD1.png
Attachment 4: QPD2.png
QPD2.png
  162   Fri Aug 30 12:22:56 2013 KojiOpticsGeneralH1 OMC Cavity side UV gluing

H1 OMC Cavity side optics was glued on the breadboard

Curved mirror gluing

- Applied the UV glues to CM1/CM2 prisms.

- Checked the spot positions on the curved mirrors

- Apply 50V to CM1

- Measure the FSR and TMS while the cavity was locked.

FSR: 261.70925MHz
TMS_V: 57.60500MHz
TMS_H: 57.94125MHz

=> Cavity round trip length of 1.1455m
=> TMS/FSR = {0.220111, 0.221395}

First accidental resonance is the lower sideband at 28th order modes.

Carrier 9th-order HOM: 2.9~7.6 line width away
Upper Sideband 13th-order HOM: 14.1-20.7 LW away
Lower Sideband 19th-order HOM: 3.3-13.1 LW away

- As this result was satisfactory, the UV illumination was zapped. It did not change the alignment. The cavity was kept locked during the illumination.

Peripheral optics gluing

- QPD path BS/Steering Mirrors were glued
- DCPD path BS was glued

The UV glue was applied to the optics.
Then the optics were placed on the breadboard along with the fixture.

Placed the dummy QPD/DCPD mount with the alignment disks.
The horizontal positions of the spots were well with in the horizontal range of the mounts.
 The UV illumination was zapped. Checked the alignment again and no problem was found.

ELOG V3.1.3-